Please show answers for both questions neatly and label

A sample of a substance has an initial temperature of 295.1 K. After absorbing 2913 J of heat, the temperature of the substanIf the Kb of a weak base is 6.0 x 10-M, what is the pH of a 0.33 M solution of this base? pH =
Answer #1

1)heat absorbed = mass of substance x specific heat x rise in temperature

Given

heat absorbed =2913 J

mass = ??

specific heat = 5.19J/g.K

rise in temperature = final-initial temperautre = 315.8 -295.1=20.7 K

substituting the values

2913J = mass x 5.19J/g.K x 20.7K

Thus mass of substance = 27.11 g

2)

Kb of base = 6.0x10-6

Thus pKb of base = - log kb = -log 6.0x10-6 = 5.22

molarity of base C =0.33M

The pOH of a weak base is given by

pOH = 1/2[pKb-logC]

= 1/2[ 5.22 -log 0.33] =2.85

Thus pH of solutionof base = 14 - pOH = 14-2.85 =11.15